A swimming pool can be filled by a large pipe in 4 hrs. and by a small pipe in 6 hrs. A third pipe can empty the pool in 3 hrs. How long would it take to fill the pool if all three pipes were open at the same time?
Answer by nerdybill(7384)

A swimming pool can be filled by a large pipe in 4 hrs. and by a small pipe in 6 hrs. A third pipe can empty the pool in 3 hrs. How long would it take to fill the pool if all three pipes were open at the same time?
.
Let x = time (hours) for all three pipe to fill pool
then
x(1/4 + 1/6 + 1/3) = 1
Multiply both sides by 12:
x(3 + 2 + 4) = 12
x(9) = 12
x = 12/9
x = 4/3
x = 1 and 1/3 hours
or
x = 1 hour and 20 minutes